OpenStudy (anonymous):

standard form ; x-intercept 4 and y-intercept -3

OpenStudy (amistre64):

well, the slope then is simply y/x if that helps

OpenStudy (amistre64):

might wanna change the sign tho :)

OpenStudy (anonymous):

so 4/-3 or 4*-3 ?

OpenStudy (amistre64):

-(-3/4) the change in y is: +3 ---- ; so slope is 3/4 the change in x is: +4

OpenStudy (amistre64):

we can use this info to construct the slope intercept form: y = mx + b y = (3/4)x -3 and algebra it into other forms as needed

OpenStudy (anonymous):

the formula is (x / x-intercept) + (y / y-intercept) = 1

OpenStudy (amistre64):

yes, that would amount to the intercept form :)

OpenStudy (anonymous):

u can use the above formula directly when given the intercepts

OpenStudy (amistre64):

i havent seen to many people that are aware if the form tho; kudos

OpenStudy (anonymous):

what would we do to get the graphing form of it?